OpenStudy (anonymous):

I am terrible at graphs and need to create a graph that shows the slope of a straight line use the point–slope form, y - y1 = m(x- x1.) The gap is 8-7-8 therefore the next gap should be 7 So 51+7 = 58

OpenStudy (anonymous):

Point slope form requires coordinates. Do you have any ordered pairs?

OpenStudy (anonymous):

Sales of 16-Inch Rims Year 1996 1998 2000 2002 Percentage 28 36 43 51

OpenStudy (anonymous):

(1998-1996)/(36-28) slope is 2/8 Ok. Assuming "year" is the "Y" axis- We know M is 2/8, or 1/4. Let (28,1996) Represent the first pair. (Y-1996)=1/4(X-28) Y-1996=1/4X - 28 y= 1/4x+1968

OpenStudy (anonymous):

sorry... y-1996=1/4x-7 y=1/4x+1989

OpenStudy (anonymous):

so this is all there is to creating a slope of a straight graph? so, if I wanted to created a actual graph I would go to word document and select the line graph and then what do I enter for the x and y to created the actual graph. Sorry I have never had to create one before!

OpenStudy (anonymous):

Your Y-intercept is 1989, and you rise 1, run 4 to create your line.

OpenStudy (anonymous):

Thanks so much for your help!